The correct answer is option 3 i.e. 2529.
Given that
P(A∪B)=95,P(Aˉ∪Bˉ)=2713,P(A)=1811
Formula used
P(Aˉ∪Bˉ) = 1 - P(A∩B)
P(A∪B)=P(A)+P(B)−P(A∩B)
Substituting the values we get
P(A∩B) = 1 - 2713 = 2714
Hence, P(B) = 95−1811+2714 = 5425
Hence, P(Bˉ) = 1 - 5425 =5429
Hence, the odds against event B = P(Bˉ) / P(B) = 2529.